(5)(-3)(9)(0)(-6)(7)
Found 2 solutions by scoop, Prithwis:
Answer by scoop(27)

Whenever you multiply anything by 0, the answer will always be 0
5*-3=-15
-15*9=-135
-135*0=0
0*-6=0
0*7=0
0 is your answer

Answer by Prithwis(166)

Since one of the numbers is zero, the product will always be zero; regardless of what the other numbers are.